OpenStudy (anonymous):

Solve each equation for y. Then find the value of y for each value of x. 1. y+2x=5; x=-1, 0, 3 2. 5x=-4y+4; x=1, 2, 3

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

please help me

OpenStudy (texaschic101):

1. y + 2x = 5 y = -2x + 5 <-- solved for y when x = -1 when x = 0 when x = 3 y = -2(-1) + 5 y = -2(0) + 5 y = -2(3) + 5 y = 2 + 5 y = 0 + 5 y = -6 + 5 y = 7 y = 5 y = -1 you understand how I did that one ??

OpenStudy (anonymous):

yeah. you just plugged in the numbers for x to find y

OpenStudy (texaschic101):

exactly....but it is first asking you to solve for y.....although, even if you didn't solve for y, you could still sub in the x value into the original equation to find y

OpenStudy (texaschic101):

5x = -4y + 4 4y = -5x + 4 <-- this is solved for y now just sub in your x points....and find a numerical value for y

OpenStudy (texaschic101):

wait..messed up.. 4y = -5x + 4 y = -5/4x + 1 <--- this is the right one...sorry about that

OpenStudy (texaschic101):

y = -5/4x + 1 when x = 1 when x = 2 when x = 3 y = -5/4(1) + 1 y = -5/4(2) + 1 y = -5/4(3)+ 1 y = -5/4 + 1 y = -10/4 + 1 y = -15/4 + 1 y = -5/4 + 4/4 y = -10/4 + 4/4 y = -15/4 + 4/4 y = -1/4 y = -6/4 or - 3/2 y = - 11/4

OpenStudy (anonymous):

thanks

OpenStudy (texaschic101):

sure thing :)